OpenStudy (kj4uts):

Which of the following are right triangles? Check all that apply. A. A triangle with side lengths 6 inches, 8 inches, 10 inches B. A triangle with side lengths 8, 15, 17 C. A triangle with side lengths 5, 12, 13 D. A triangle with side lengths 4, 5, 6 Please explain answer. Is there a way to solve this or ratios to remember? Thank you!

jimthompson5910 (jim_thompson5910):

Given a triangle with side lengths a,b,c. The length c is the longest side If you can show that a^2 + b^2 = c^2 is true, then the triangle is a right triangle. This is the converse of the pythagorean theorem.

jimthompson5910 (jim_thompson5910):

Example a = 3 b = 4 c = 5 a^2 + b^2 = c^2 3^2 + 4^2 = 5^2 9 + 16 = 25 25 = 25 the triangle with sides 3,4,5 is a right triangle